OpenStudy (anonymous):

system of equations: 2x-1/3y=6 3x-5/6y=-13 (I am in algebra just in case someone asks)

OpenStudy (anonymous):

x=3.8148 y=4.8888

OpenStudy (anonymous):

i believe i need it in fraction....

OpenStudy (anonymous):

can you show me how to do the work?

OpenStudy (anonymous):

sorry the above answer is incorrect.............. the correct sol: the eqs can be rewritten as: 6x-y=18.................i 18x-5y=-78............ii eq (i) X 3 gives 18x-3y=54...........iii subtract eq(ii) from eq(iii) gives; 2y= 132 or, y=66 put this y in eq(i) gives x= 14.......ans